# Tool-as-Interface: Learning Robot Policies from Human Tool Usage through Imitation Learning

### v1

@@ Abstract
Tool use is critical for enabling robots to perform complex real-world tasks, and leveraging human tool-use data can be instrumental for teaching robots. However, existing data collection methods like teleoperation are slow, prone to control delays, and unsuitable for dynamic tasks. In contrast, human natural data—where humans directly perform tasks with tools—offers natural, unstructured interactions that are both efficient and easy to collect. Building on the insight that humans and robots can share the same tools, we propose a framework to transfer tool-use knowledge from human data to robots. Using two RGB cameras, our method generates 3D reconstruction, applies Gaussian splatting for novel view augmentation, employs segmentation models to extract embodiment-agnostic observations, and leverages task-space tool-action representations to train visuomotor policies. We validate our approach on diverse real-world tasks, including meatball scooping, pan flipping, wine bottle balancing, and other complex tasks. Our method achieves a 71% higher average success rate compared to diffusion policies trained with teleoperation data and reduces data collection time by 77%, with some tasks solvable only by our framework. Compared to hand-held gripper, UMI [16], our method cuts data collection time by 41%. Additionally, our method bridges the embodiment gap, improves robustness to variations in camera viewpoints and robot configurations, and generalizes effectively across objects and spatial setups.
@@
I Introduction
Tool use is essential to how humans interact with and transform their environment [67, 41]. For instance, humans use a pan to fry food and flip it, ensuring even cooking on both sides. Despite its significance, tool use beyond parallel jaw grippers remains underexplored in robotics, with research primarily focused on simpler tasks like grasping and pick-and-place operations [6, 46, 10, 36, 35]. In this paper, we focus on cost-effective data collection and efficient training of robot policies to rapidly acquire tool-use skills.
Imitation learning provides a promising pathway for robots to acquire tool-use skills by directly learning from human demonstrations [21, 26, 27, 29]. The paradigm excels in handling diverse tool-use tasks, as it bypasses the need for task-specific programming by relying solely on human demonstrations. However, its full potential hinges on addressing key challenges in collecting high-quality training data. Various teleoperation systems [33, 56, 8, 20, 29, 50, 84, 68, 39, 66, 45, 9] and hand-held grippers [65, 19, 53, 47, 43] have been developed to facilitate the collection of high-quality data. Teleoperation methods, such as kinematic replication and hand or body retargeting, show great potential [85, 22, 74, 62, 48]. However, their reliance on direct access to robot hardware limits both practicality and scalability. Hand-held grippers [57, 16] offer an alternative by enabling demonstrations in diverse environments. While they reduce dependency on robotic systems, their high costs and the technical expertise required for tasks like 3D printing and assembly restrict their accessibility to a specialized group of users.
To address these limitations, we turn to human interaction data — a natural, intuitive method through which humans interact with their environment during everyday activities without relying on external devices or specialized setups. Human interaction data refers to the natural process in which humans use their hands to operate tools and interact with and manipulate the environment freely. Unlike controlled demonstrations that require expensive hardware or meticulous preparation, human interaction data involves the spontaneous use of tools to interact with the environment. Human interaction data is an accessible, scalable, and cost-effective approach to data collection, requiring no prior knowledge or technical expertise, such as 3D printing or assembly, from participants. However, existing methods struggle to fully harness the potential of human interaction data. Key challenges include the embodiment gap and the reliance on single-view data, which limits the insights that can be drawn from interaction data [60, 4, 38, 5, 64, 42, 55].
Our framework addresses these challenges by leveraging the observation that humans and robots can share the same tools. We propose a novel approach that utilizes human interaction data to train robust and adaptable robot policies for diverse tool-use tasks (Figure 1). Our method minimizes reliance on expensive hardware, making data collection more scalable and accessible to non-experts. By capturing 3D information using two RGB cameras and generating 3D reconstructions, our method enables view-invariant policy learning through novel view augmentation. To facilitate direct policy transfer from human interaction data to robotic systems, we employ a segmentation model to filter out embodiment-specific information. Additionally, we leverage task-space tool-action representations to ensure robustness to variations in robot base configurations.
Our contributions are as follows:
  1. 1.
    We propose a novel framework that leverages two-view human interaction data to enable scalable, intuitive, and cost-effective data collection for training robot policies on complex tool-use tasks.
  2. 2.
    We validate the effectiveness of our approach on diverse challenging real-world tool-use tasks (e.g., nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking). Our method achieves a 71% higher success and 77% less data collection time than diffusion policies trained on SpaceMouse [17] or Gello [74] data, with some tasks solvable only by our method. Our method also outperforms handheld grippers like UMI [16], reducing collection time by 41%.
  3. 3.
    We provide an extensive analysis of our method’s robustness under varying conditions, including changes in camera poses, robot base movements, and human-induced perturbations. Additionally, we conduct ablation studies to evaluate the effects of different design choices on policy performance, including embodiment segmentation, random cropping, and novel view augmentation.
II Related Works
Figure 2: Policy Design. Human interaction data was collected using two RGB cameras and processed through the foundation model MASt3R [34] to generate 3D reconstructions. Using 3D Gaussian splatting, we sampled novel views to augument the dataset. The human hand (embodiment) was segmented out from the images to create embodiment-agnostic observations, which serve as inputs to the policy. To label actions for policy training, a pose estimation model (FoundationPose [71] in this work) was used to extract the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$. A coordinate transformation was then applied to compute the tool’s pose in the task space, $T_{\text{tool}}^{\text{task}}$. Finally, a diffusion policy was implemented as the visuomotor policy to enable effective learning and execution.
II-A Data Collection for Robot Learning
@@
High-quality data is essential for training robots to learn and generalize across tasks. Simulation data has been widely used for its cost-effectiveness and scalability [86, 44, 28, 81], but the gap between simulation and real-world performance remains a persistent challenge. To overcome this, many researchers have turned to real-world teleoperated demonstrations, which offer more reliable transferability by minimizing the domain gap between training and testing environments [88, 14, 15, 56, 73]. Advancements in leader-follower devices, like ALOHA [85, 22] and GELLO [74], have simplified robot demonstration data collection but remain tied to specific robot platforms. More recently, portable tools such as hand-held grippers, e.g., UMI [16] and LEGATO [57], have emerged as a promising alternative for in-the-wild data collection. Yet, their high cost and the need for custom robot modifications continue to limit widespread adoption. Tool-based policy representations have emerged as an effective way to collect data for robot learning. MimicTouch uses tactile-based tools for contact-rich manipulation [79], while ScrewMimic models bimanual tasks as constrained screw motions for learning from human videos [3]. However, tactile methods require extra hardware, and the screw motion assumption may not hold. Another line of work by Wen et al. [70] aims to learn category-level representations from a single demo to transfer pose trajectories across similar objects [70], but assumes a static target (e.g., battery slot), limiting real-world applicability in real-world tasks where the spatial configuration of objects may change. Unlike these methods, our method only requires natural human data, without assuming access to tactile sensors or requiring constrained screw motion models, making it significantly more cost-effective, scalable, and accessible.
II-B Cross-Embodiment Policy Learning
@@
Cross-embodiment policy learning enables robots to transfer policies across embodiments, such as those with different kinematic structures [25, 52, 77]. Prior work has explored conditioning policies on embodiment representations using multi-embodiment datasets [18, 69, 80, 13, 24, 52, 2], but these approaches often face challenges in effectively leveraging human interaction data. Recent approaches utilize human data, such as estimating point flow from human video [72] or generating latent plan or high-level plans from human data [37, 68]. However, their dependence on robot data for low-level control limits scalability when robot data is expensive or difficult to collect. Additionally, prior works highlight the importance of masking human and robot embodiments for visual consistency [4, 31]. However, Bahl et al. [4] relies on predefined motion primitives, while Kareer et al. [31] still requires robot data with human data as augmentation. Our approach adopts a similar masking idea but enables robots to learn freely, even agile motions from human videos, without any robot data. Another line of work tokenizes observation inputs and action outputs into a unified transformer network, enabling generalized policy learning across embodiments [59, 76]. However, these methods require large models and extensive datasets, making them resource-intensive and time-consuming to train, and lack the capability for direct policy transfer between embodiments. Overall, reliance on robot-specific hardware and data restricts scalability and accessibility. In contrast, our method leverages natural human data, eliminating the need for robot data as a training source.
II-C Cross-Viewpoint Policy Learning
@@
Viewpoints can change as robots manipulate or interact with their environment. Existing approaches address viewpoint changes by learning view-invariant latent representations [58, 11], leveraging equivariant 3D representations [63, 23, 89], or augmenting input data with diverse viewpoints [51, 87]. An early attempt by Sharma et al. [61] introduced a hierarchical model that generates first-person sub-goals from third-person demonstrations via a high-level module, while a low-level embodiment-specific module predicts actions. Recent work by Chen et al. [12] leveraged the SAM model [32] and ControlNet [83] to transform the viewpoint of one robot into another, performing view synthesis with ZeroNVS [54] to augment datasets and enhance policy learning across robots and viewpoints. Another approach by Yuan et al. [82] combined reinforcement learning with multi-view representation learning and a Spatial Transformer Network (STN) to improve policy robustness in visually complex environments. However, these methods face limitations such as scalability issues in hierarchical models, the complexity of controlled image editing, high computational costs in view synthesis, and the need to build sophisticated simulators for reinforcement learning when adapting to diverse environments. Compared to previous work, our method enables efficient and scalable cross-viewpoint transfer from human data by leveraging two RGB images with the MASt3R [34] foundation model for 3D reconstruction and employing Gaussian splatting for fast and cost-effective novel view synthesis.
III Problem Statement
@@
We formulate the robotic manipulation task as a Markov Decision Process, where the goal is to learn a policy $\pi:\mathcal{O}^{r}\to\mathcal{A}$ that enables a robot to perform a given task. The robot’s observation space, $\mathcal{O}^{r}$, consists of a single-view RGB image $I^{r}\in\mathbb{R}^{128\times 128\times 3}$ and proprioception data $x^{r}\in SE(3)$. To train the policy, we use an imitation dataset composed of $N$ human play, denoted as $D=\{(O_{0}^{h},O_{1}^{h},\dots)\}_{n=1}^{N}$. Each human play observation $\mathcal{O}^{h}$ consists of two RGB images captured from different viewpoints: $\mathcal{O}^{h}=\{I^{v1},I^{v2}\}$, where $I^{vi}\in\mathbb{R}^{480\times 640\times 3}$. We preprocess the dataset to get the action by using a 6D pose estimation and tracking model, resulting in $D=\{(O_{0}^{h},a_{0},O_{1}^{h},a_{1},\dots)\}_{n=1}^{N}$, where each action $a\in SE(3)$.
Our objective is to train a robot to perform the same task as demonstrated in the human dataset while overcoming the embodiment differences. We assume that the tool is rigid, the tool and the end effector are rigidly attached during deployment, and the transformation between the tool and the end effector is estimated once before deployment. Our approach aims to allows the robot generalize task execution across embodiments while retaining the core skills demonstrated by humans.
IV Our Robot Policy Learning Framework

### v2

# Tool-as-Interface: Learning Robot Policies from Observing Human Tool Use

@@ Abstract
Tool use is essential for enabling robots to perform complex real-world tasks, but learning such skills requires extensive datasets. While teleoperation is widely used, it is slow, delay-sensitive, and poorly suited for dynamic tasks. In contrast, human videos provide a natural way for data collection without specialized hardware, though they pose challenges on robot learning due to viewpoint variations and embodiment gaps. To address these challenges, we propose a framework that transfers tool-use knowledge from humans to robots. To improve the policy’s robustness to viewpoint variations, we use two RGB cameras to reconstruct 3D scenes and apply Gaussian splatting for novel view synthesis. We reduce the embodiment gap using segmented observations and tool-centric, task-space actions to achieve embodiment-invariant visuomotor policy learning. We demonstrate our framework’s effectiveness across a diverse suite of tool-use tasks, where our learned policy shows strong generalization and robustness to human perturbations, camera motion, and robot base movement. Our method achieves a 71% improvement in task success over teleoperation-based diffusion policies and dramatically reduces data collection time by 77% and 41% compared to teleoperation and the state-of-the-art interface, respectively.
@@
Keywords: Tool Use, Data Collection, Learning from Video
1 Introduction
Tool use enables humans to perform complex tasks by extending their physical capabilities. In contrast, robotic systems remain largely limited to grasping and pick-and-place operations [1, 2, 3, 4, 5, 6, 7]. To enable richer manipulation skills, robots must learn to use diverse tools in dynamic environments. This work focuses on the efficient training of robot policies for tool use, with an emphasis on scalable and low-cost data collection.
Imitation learning (IL) is a promising way to acquire tool-use skills from human demonstrations [8, 9, 10]. Prior work has leveraged teleoperation platforms [11, 12, 10] and hand-held grippers [13, 14] to provide precise supervision. However, these systems often require expensive hardware, 3D-printed tools, or expert calibration, hindering scalability across diverse users and environments.
As a scalable alternative, we learn from natural human manipulation videos. While abundant, this data introduces challenges from viewpoint differences and human-robot embodiment gaps [15, 16, 17]. We introduce a new framework that leverages two-view human manipulation videos and appropriate representations of states and actions to train robot policies for tool use. To close the viewpoint gap, our system uses 3D scene reconstruction and novel view synthesis to train a viewpoint-invariant policy. To bridge the embodiment mismatch, we filter out embodiment-specific features like hands using segmentation, and employ a task-space, tool-centric action representation supports robustness to robot base variation (Figure 1).
Our contributions are as follows: (1) We introduce a framework for scalable, intuitive, and cost-effective data collection for robot tool-use learning, using two-view human manipulation videos without requiring teleoperation or specialized hardware; (2) We demonstrate strong generalization across diverse real-world tool-use tasks (e.g., nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking) achieving a 71% higher success rate and 77% reduction in data collection time compared to diffusion policies trained on SpaceMouse [18] or Gello [19], and a 41% improvement over handheld grippers like UMI [14]; and (3) We provide a detailed robustness analysis, evaluating performance under changes in viewpoint, robot base configuration, and human motion, along with ablations on segmentation, novel view synthesis, and random cropping.
2 Related Works
@@
Data Collection for Robot Learning: High-quality data is essential for teaching robots new skills. Simulation offers scalability and low cost [20, 21, 22, 23], but the sim-to-real gap remains challenging. Learning from real-world teleoperated demonstrations helps train policies with minimal distribution shift between training and testing [24, 25, 26, 27, 28]. Leader-follower systems (e.g., ALOHA [29, 30], GELLO [19]) simplify teleoperation by providing intuitive kinematic replication but require real robots and are costly. Portable hand-held grippers (e.g., UMI [14], LEGATO [13]) enable flexible data collection but still require specialized hardware. Tool-based policy representations offer another direction. MimicTouch leverages tactile feedback for contact-rich skills [31], while ScrewMimic models bimanual tasks as constrained screw motions for learning from human videos [32]. However, tactile methods require extra hardware, and the screw motion assumption may not hold. Another approach transfers single-demonstration trajectories across objects but assumes static object configurations [33], which will not work with changing spatial configurations. In contrast, our method uses natural human demonstrations without tactile sensors, special tools, or motion constraints—enabling scalable, low-cost data collection in unconstrained environments.
@@
Cross-Embodiment Policy Learning: Cross-embodiment learning enables policy transfer across robots with diverse kinematic structures [34, 35, 36]. Prior approaches relying on multi-embodiment datasets [37, 38, 39, 40, 41, 35, 42] focus solely on data from different robot embodiments and struggle to leverage human demonstrations effectively. Recent methods tokenize observations and actions into unified transformer networks [43, 44], but require large models, extensive datasets, and cannot directly transfer policies across embodiments. Other efforts use human data to estimate point flow [45] or generate high-level plans [46, 11], but still depend on robot-generated data for low-level control. Additionally, prior works emphasize visual consistency through embodiment masking [16, 47]. However, Bahl et al. [16] relies on predefined motion primitives and Kareer et al. [47] requires robot data for augmentation. As a step further, our approach adopts a similar masking idea but enables robots to learn freely, even agile motions from human videos without any robot data.
@@
Cross-Viewpoint Policy Learning: Robots can encounter viewpoint variations when interacting with their environments. Prior work addresses this by learning view-invariant representations [48, 49], equivariant 3D features [50, 51, 52], or augmenting datasets with varied viewpoints [53, 54]. An early attempt by Sharma et al. [55] introduced a hierarchical model mapping third-person demonstrations to first-person sub-goals, with a low-level module predicting actions. Recent work by Chen et al. [56] leverages SAM [57] and ControlNet [58] to transform robot images across different robots, employing ZeroNVS [59] for novel-view synthesis to augment datasets and improve policy generalization across viewpoints. Another approach by Yuan et al. [60] integrates reinforcement learning, multi-view representation learning, and a Spatial Transformer Network to enhance policy robustness in visually complex environments. Despite promising results, these methods face scalability issues, reliance on computationally intensive view synthesis, or the need for tailored simulators. In contrast, our approach achieves efficient and scalable cross-viewpoint transfer directly from human data by leveraging the MASt3R [61] foundation model for 3D reconstruction from two RGB images and employing Gaussian splatting for fast, cost-effective novel-view synthesis.
3 Tool-as-Interface Framework
Figure 2: Policy Design. Human manipulation data was collected using two cameras and processed through the foundation model MASt3R [61] to generate 3D reconstructions. Using 3D Gaussian splatting, we sampled novel views to augment the dataset. Human hands were segmented to create embodiment-agnostic observations as policy inputs. For action labeling, FoundationPose [62] estimated the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$, which was transformed into task space, $T_{\text{tool}}^{\text{task}}$. A diffusion model was then trained as the visuomotor policy.
@@ 3 Tool-as-Interface Framework
Problem Statement: We formulate robotic manipulation as a Markov Decision Process (MDP), where the goal is to learn a policy $\pi:\mathcal{O}^{r}\to\mathcal{A}$ that enables a robot to perform a given task. The robot’s observation space $\mathcal{O}^{r}$ consists of a single-view RGB image $I^{r}\in\mathcal{I}^{r}$ and proprioceptive data $x^{r}\in SE(3)$, where each $I^{r}$ is a tensor in $\mathbb{R}^{128\times 128\times 3}$. We train the policy using an imitation dataset of $N$ human demonstrations, $D={(\mathcal{O}_{0}^{h},\mathcal{O}_{1}^{h},\dots)}_{n=1}^{N}$, where each $\mathcal{O}^{h}=\{I_{v1}^{h},I_{v2}^{h}\}$ contains two RGB images captured from different viewpoints and each $I_{vi}^{h}\in\mathcal{I}^{h}$ is a tensor in $\mathbb{R}^{480\times 640\times 3}$. We preprocess the dataset to infer actions using a 6D pose estimation and tracking model, resulting in $D=\{(\mathcal{O}_{0}^{h},a_{0},\mathcal{O}_{1}^{h},a_{1},\dots)\}_{n=1}^{N}$, where each action $a\in SE(3)$. To bridge the embodiment gap between humans and robots, we assume the tool is rigidly attached to both the human hand (implicitly) and the robot end-effector (explicitly), with a fixed transformation estimated prior to deployment. Under this setup, the robot can reproduce human-demonstrated tool trajectories, enabling policy transfer across embodiments while preserving task-relevant behaviors (Figure 2).
Tool-Centric Demonstrations for Robot Manipulation: We leverage the fact that both humans and robots can operate the same physical tools to facilitate policy learning. Tools serve as a shared interface for interacting with objects, enabling the direct transfer of human demonstrations with minimal embodiment-specific adaptation. Unlike prior work focused on grasping or pick-and-place tasks [14, 13, 63], our approach enables robots to perform complex interactions using everyday tools. Our formulation abstracts actions to the tool pose, reducing morphological dependence and promoting policy generalization across embodiments. It also simplifies data collection by eliminating the need for robot-specific demonstrations. Humans can naturally manipulate tools by hand without extra instrumentation. For deployment, robots either rigidly grasp the tool, as demonstrated with a Kinova Gen3 arm, or attach it using a custom fast tool changer described in Appendix B.2.1 and shown in Figure 10, compatible with ISO 9409-1-50-4-M6 flanges.
Perception Alignment Across Embodiments: To enable cross-embodiment policy transfer, we align human and robot observations within a shared visual space $\mathcal{I}^{s}$ by applying a feature extraction function $g:\mathcal{I}^{h}\cup\mathcal{I}^{r}\to\mathcal{I}^{s}$. We instantiate $g$ with Grounded-SAM [64], using prompts such as “human hand” and “robot arm” to mask out embodiment-specific regions—human hands during training and robotic arms during deployment. Masking out these regions ensures that only task-relevant visual information (e.g., tools and objects) remains visible in both phases. By minimizing visual discrepancies between human and robot data, the feature extraction process reduces embodiment-specific bias and improves generalization across embodiments.
3D-Aware View Augmentation: We use cameras for data collection due to their widespread availability—over 7.14 billion smartphones are equipped with them [65]. However, single-camera setups suffer from scale ambiguity and limited 3D perception and are sensitive to viewpoint changes.
3D Reconstruction: To address this, we use MASt3R [61], an image-matching model that reconstructs accurate 3D environments from just two RGB images—eliminating the need for depth sensors, which are less common and more power-hungry. Two cameras suffice to avoid scale ambiguity inherent in monocular settings. MASt3R produces high-quality point clouds without requiring known camera extrinsics or intrinsics by globally aligning multi-view features.
View Synthesis and Augmentation: 3D Gaussian splatting synthesizes novel viewpoints from the reconstructed scene, allowing the robot to observe interactions from multiple angles—even when only two views are available. The resulting perspectives augment the training data, increasing visual diversity and improving policy learning. To further enhance robustness and generalization, random cropping is applied, following diffusion policy [25, 26].
Tool-Centric Action Representation and Policy Deployment: To support general tool usage, we propose a task-frame, tool-centric action representation denoted as $T_{\text{tool}}^{\text{task}}$, which describes the tool’s motion independently of human or robot morphology, camera pose, or base configuration. This invariant formulation enables robust policy transfer across different embodiments and viewpoints. As shown in Figure 3, the tool’s pose is first estimated in the camera frame using a 6D pose estimation model (e.g., FoundationPose [62]) as $T_{\text{tool}}^{\text{camera}}$, and then transformed into the task frame:
where $T_{\text{camera}}^{\text{task}}$ denotes the transformation from the camera to the task frame.
Figure 3: Coordinate System Diagram. The diagram shows the Camera, Tool, Task space, End-Effector (EEF), and Base frames. The action is represented as $T_{\text{tool}}^{\text{task}}$.
A diffusion policy [25] maps a single-view RGB image to a predicted SE(3) action $T_{\text{tool}}^{\text{task}}$. At deployment, the robot command is computed by converting the prediction to the end-effector frame. For stationary robots, the task frame aligns with the base frame; for mobile platforms, base movement is compensated using $T_{\text{task}}^{\text{base}}$. The resulting end-effector pose is given by:
where $T_{\text{eef}}^{\text{tool}}$ is the known fixed transform between the tool and the robot end-effector.
4 Policy Evaluations
Our evaluations aim to assess our framework across three dimensions: reliability (how consistently and successfully the learned policies perform), execution efficiency (how smooth and natural the resulting behaviors are), and versatility (how well the framework adapts to diverse tasks and generalizes across conditions).
Experimental Tasks Overview: We evaluate five real-world robotic tasks on Kinova Gen3 and UR5e robots, involving precision manipulation, dynamic object handling, and dexterous tool use. Policies use RGB inputs from RealSense D415 cameras and handle variations in object positions and camera poses. Tasks include: (1) Nail Hammering – Precise striking of a small target, (2) Meatball Scooping – Contact-sensitive rolling object manipulation, (3) Pan Flipping – Fast, dynamic flipping with varied objects, (4) Wine Balancing – Gravity-aware placement into an unstable rack, and (5) Soccer Ball Kicking – Dynamic interception and obstacle avoidance. Full details in Appendix B.1.
Baselines: We evaluate the effectiveness and efficiency of learning directly from human manipulation videos without relying on robot-generated data. We benchmark against a diffusion policy trained on robot demonstrations and UMI [14], a hand-held gripper method. Robot demonstrations are collected using SpaceMouse or Gello [19] under identical time budgets. Additionally, we conduct ablations to assess random cropping before policy training, novel view synthesis data augmentation, and embodiment segmentation. To further illustrate the advantages of our approach, we compare trajectory rollouts for a meatball-scooping episode, highlighting how our method is more sample-efficient and less prone to distribution shifts by eliminating excessive waypoints.
5 Experiment Results
Capabilities and Effectiveness: Table 1 summarizes our real-world results, showing that our framework consistently achieves higher success rates across all tasks compared to baselines. We also compare against the stronger hand-held gripper baseline UMI [14] (Table 2). In our default setup, SLAM-based mapping failed due to low environmental texture, so we added a textured background to support reliable mapping for UMI. For the nail hammering task, we evaluated UMI with 25 demonstrations (matching our collection time) and 100 demonstrations (to assess ideal performance). UMI fails all 13 trials with 25 demonstrations but succeeded with 100. It was also inapplicable
Table 1: Task Success Rates and Completion Times. Success rates show completed trials over total. “DP” is trained on teleoperation; “Not Feasible” means failure.
Task Method Success Time (s)
Hammer DP 0/13
Ours 13/13 11.0
Scoop DP 5/12 42.0
Ours 10/12 12.4
Pan: Egg DP Not Feasible
Ours 12/12 1.5
Pan: Bun DP Not Feasible
Ours 9/12 1.9
Pan: Patty DP Not Feasible
Ours 10/12 2.3
Wine Balance DP Not Feasible
Ours 8/10 30.9
Soccer Kick DP Not Feasible
Ours 6/10 2.0
to wine balancing and pan flipping due to contact and inertial challenges, and struggled in soccer kicking due to localization failures. In contrast, our method demonstrates reliable performance across all tasks: accurately detecting spatial locations (nail hammering, meatball scooping), performing high-speed motions (pan flipping), precisely inserting wine bottles, and swiftly reacting in soccer kicking. This strong performance is enabled by collecting significantly larger and more diverse episodes within the same data collection timeframe (see Section 6), enabling robust policy training. Our approach overcomes limitations of teleoperation tools like Gello and SpaceMouse, enabling data collection for scenarios they struggle to handle. Qualitative policy rollouts are shown in Figure 9 in Appendix.
Policy Execution Efficiency: Our framework demonstrates high execution efficiency, achieving shorter task completion times and smoother action trajectories than baselines (Table 1). This efficiency stems from the natural human data, which captures the fluidity and speed of real-world activities, resulting in higher-quality training trajectories. In contrast, teleoperated demonstrations
Table 2: Task success rates comparing our method with the hand-held gripper-based method on Nail Hammering.
Method Demo Time & Count Success
UMI [14] 180s (25) 0/13
UMI 720s (100) 13/13
Ours 180s (40) 13/13
often produce slower, less natural motions, limiting dynamic performance. By leveraging more realistic data, our framework accelerates execution while enhancing motion quality, making it suited for real-world applications. See Appendix C for a detailed comparison of policy rollouts.
Benefits of Tool-Based Action Representation in Task Space: Using the tool pose in the camera frame works with a static camera but fails under camera movement due to unreliable real-time tracking and incorrect end-effector positioning in the base frame. Similarly, representing actions in the base frame fails under base movement due to the assumption of a fixed base-to-workspace transform. In contrast, representing actions in task space is invariant to both camera and base movement, enabling robust execution even under large viewpoint shifts and base movements.
Effects of Random Cropping and View Augmentation: Our experiments show that random cropping (RC) and view augmentation (VA) together enhance policy robustness to camera pose variations. RC improves resilience to minor perturbations such as small movements or shaking, while VA exposes the model to a broader distribution of viewpoints during training. We evaluated these techniques on the nail hammering task (Figure 4), comparing three models: one trained with both RC and VA, one with RC only, and one without either. The combined use of RC and VA significantly expands the range of camera configurations under which the policy can successfully operate.
Generalization:
Spatial Generalization: We evaluated spatial generalization by varying initial conditions across tasks: nail positions for hammering, meatball locations for scooping, goalkeeper setups for soccer ball kicking, and object poses across the pan for flipping (illustrated in Figure 11 in Appendix).
Object Generalization: Our method generalizes effectively to different objects in the pan-flipping task, including the egg and burger bun seen during training, and a 3D-printed meat patty (illustrated in Figure 11, second column, in Appendix). The policy learns to tilt the pan to slide the object into a corner, then flick it to achieve a successful flip, enabling robust generalization across object types.
Tool Generalization: We evaluated tool generalization by testing the policy with five different pans: large, medium, small, tiny, and square. The policy was trained using demonstrations with the large, medium, and square pans and evaluated on all five, with 12 trials per pan under varying initial configurations (illustrated in Figure 13 in Appendix). It achieved high success rates on the trained pans (large and medium). Performance declined on smaller pans due to limited surface area, and on the square pan due to shallow edges causing the bun to slide out during flipping.
Figure 6: Human Perturbation Robustness. The robot handles human-induced perturbations across three tasks: (1) In nail hammering, it tracked a manually moved nail; (2) In meatball scooping, it located and scooped new meatballs thrown mid-task; and (3) In egg flipping, it recovered the egg after human repositioning.
Robustness:
Camera Pose Robustness:
We evaluated the policy’s ability to handle camera pose variations by introducing camera shaking in three tasks: meatball scooping, nail hammering, and pan flipping (Figure 5(a)). The first row shows the camera view, and the second shows the scene overview and shaking motion. Despite disturbances, the policy consistently completed all tasks, enabled by random cropping during training, improving adaptation to partial views and minor visual changes.
Robot Base Robustness: To assess robustness to base movement, we manually shook the robot base during execution (Figure 5(b)). When the shaking frequency exceeded the control frequency, the end effector oscillated with the base; however, the task-space action design enabled compensation and successful task completion. As shown in Figure 5(d), the policy also maintained effectiveness under simultaneous camera and base shaking.
Chicken Head Stabilization: At lower shaking frequencies, where the perturbation was slower than the robot’s control loop, the end effector exhibited a stabilization behavior similar to a chicken’s head [66] (Figure 5(c)), maintaining steady control during mild base movements.
Human Perturbation Robustness: We evaluated resilience to human interventions (Figure 6). The robot tracked moving nails, adapted to new meatballs thrown in mid-task, and re-flipped repositioned eggs, demonstrating robustness to real-time disturbances.
6 Data Collection Efficiency and Affordability
We compare data collection methods for robot imitation learning across throughput, reliability, cost, usability, and precision. Full quantitative and qualitative analysis is provided in Appendix D.
Data Collection Efficiency: Leveraging the natural dexterity and intuitive control of human hands (Figure 7(a)), our method achieves significantly higher data collection throughput compared to traditional methods. Quantitatively, human demonstrations reduce data collection time by 73% for nail hammering and 81% for meatball scooping relative to teleoperation methods (Gello, Spacemouse), as shown in Figure 8. Compared to handheld grippers (e.g., UMI [14]), human demonstrations are 41% faster in nail hammering, and further succeed in dynamic tasks where both teleoperation and handheld methods consistently fail. This efficiency enables high-throughput, low-variance data collection critical for scalable robot learning.
Reliability: In contrast, Figures 7(b) and (c) highlight typical failures for Gello, Spacemouse, and handheld grippers, including collisions, tool slippage, safety stops, and failures in dynamic or high-precision tasks. Teleoperation tools frequently suffer from latency, lack of tactile feedback, and difficulty handling rapid motions, leading to inconsistent trajectories and poor-quality demonstrations. Handheld grippers are prone to inertial slippage and loss of tool control under high forces, further limiting their applicability.
Figure 8: Quantitative Comparison of Data Collection Methods. Human hands reduce data collection time by 73% for nail hammering and 81% for meatball scooping, while maintaining low variation. Teleoperation fails in dynamic and high-precision tasks. In nail hammering, human hands are 41% faster than UMI [14], which also struggles with dynamic and low-texture environments.
Cost and Usability: Our method requires no specialized hardware, incurring zero additional cost compared to teleoperation devices, handheld grippers, and VR/AR equipment, all of which demand substantial investment and maintenance. A detailed discussion is provided in Appendix D, with a direct comparison summarized in Table 4. The intuitive nature of human demonstrations further lowers the expertise barrier, making large-scale, high-quality data collection accessible without costly infrastructure.
By eliminating hardware dependence, minimizing operational failures, and dramatically improving throughput and precision, our human-centric data collection framework enables scalable, efficient, and reliable robot imitation learning.
7 Conclusion
In this work, we presented a framework that enables robust policy training for diverse tool-use tasks by learning from human manipulation videos and bridging the embodiment gap they introduce. Unlike traditional data collection methods, which are costly and hardware-dependent, our approach democratizes data collection by eliminating the need for specialized equipment or technical expertise, making large-scale robot learning accessible and scalable. We validated the framework across challenging tasks, including nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking, demonstrating superior performance, robustness to variations in camera poses and base movements, and adaptability across 6-DOF and 7-DOF robots. By improving accessibility, scalability, and reliability, our work lays a strong foundation for advancing robotic manipulation in complex, real-world scenarios.
8 Limitations and Future Work
Acknowledgments
We thank Professor Wenzhen Yuan, Xiaoyu Zhang, Yucheng Mo, Yilong Niu, Hyoungju Lim, and Amin Mirzaee for their support and assistance in reproducing the UMI. We thank Professor Saurabh Gupta, Professor Junyi Geng, Lujie Yang, Neeloy Chakraborty, Hongkai Dai, Jacob Wagner, Shaoxiong Yao, Wei-Cheng Huang for their insightful feedback and suggestions. This work was supported by ZJU-UIUC Joint Research Center Project No. DREMES 202003, funded by Zhejiang University. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. Additionally, this work used NCSA Delta GPU at NCSA through allocation CIS240753 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The robot experiments were carried out in part in the Center for Autonomy Robotics Laboratories at the University of Illinois Urbana-Champaign.
APPENDIX
Appendix A Design Choice
A.1 Key Capabilities and Practical Benefits